Assume that the average annual salary for a worker in the United States is $31,000 and that the annual salaries for Americans are normally distributed with a standard deviation equal to $7,500. Find the following:

z(45000) = (45000-31000)/7500 = 1.8667
P(x > 45000) = P(z > 1.8667) = normalcdf(1.8667,100) = 0.0310
(B)What percentage of Americans earn above $45,000?

Answer by jello3211(7)


(A)What percentage of Americans earn below $20,000?
A)What percentage of Americans earn below $20,000?
z(20000) = (20000-31000)/7500 = -1.4667
P((x < 20,000) = P(z < -1.4667) = normalcdf(-100,-1.4667) = 0.0712
(B)What percentage of Americans earn above $45,000